OpenStudy (jordan123321):

What is the median of the data? 12, 17, 16, 10, 20, 13, 14, 14, 12, 12, 19, 18 A. 16 B. 15 C. 17 D. 14

OpenStudy (jordan123321):

@PizzaLover123

OpenStudy (er.mohd.amir):

arrange data in increasing order if number of data is n and odd then median =(n+1)/2 th term

OpenStudy (anonymous):

Median - Middle value when even if its not even you add the 2 middle numbers and divide by 2.

OpenStudy (jordan123321):

im confused

OpenStudy (anonymous):

14 + 13 = ? \[? \div 2 = \] Your answer.

OpenStudy (anonymous):

Get it? @jordan123321

OpenStudy (jordan123321):

26 / 2

OpenStudy (jordan123321):

so 13

OpenStudy (anonymous):

14 + 13 = 27 27 / 2 = 13.5 aka 14

OpenStudy (anonymous):

So your answer is ? ^.^

OpenStudy (anonymous):

Thanks for the badges ^.^